# Welfare Analysis in Dynamic Models

### v4

@@ 6.1. Average Welfare and Related Averages
Algorithm 1 Step 1 partitions the data into $L$ folds to enable sample splitting. Steps 2 and 3 estimate the value function and the dynamic dual representation $\alpha$ on each fold by minimizing cross-fit loss functions. Step 4 estimates the per-period utility $\zeta(\cdot)$ and the bias-correction terms using observations not in fold $\ell$. Step 5 computes the welfare estimate $\widehat{\delta}$ as a bias-corrected plug-in estimator. Step 6 computes the standard error of $\widehat{\delta}$ using the empirical variance of the influence function $\psi$, defined in equation (6.6).
@@ 7.4. Discussion and Related Results
In this Section, we discuss related results. Remark 7.3 discusses cross-fitting. Remark 7.4 introduces a neural network estimator of the value function. Remark 7.5 introduces a neural network estimator of the dynamic dual representation. Remark 7.6 describes the automatic property of the dynamic dual criterion function. Remark 7.7 verifies rate conditions for asymptotic theory. Remark 7.8 sketches a Dantzig estimator of the value function.
@@
Remark 7.8 (Dantzig estimator of the value function).
Earlier work has focused on an instrumental variable (IV) approach corresponding to a minimum distance problem
$\mathrm{E}[b(X)(b(X)-\beta b(X_{+}))^{\prime}]\rho_{V}-\mathrm{E}[b(X)\zeta_{0}(X)]$ (7.28)
$=G\rho_{V}-M=0.$
Unlike $G_{V}$ in (7.9), the matrix $G=\mathrm{E}[b(X)(b(X)-\beta b(X_{+}))^{\prime}]$ is not symmetric and does not correspond to a FOC of a least squares criterion function. Yet, a sparsity restriction of $\rho_{V}$ can be imposed using a Dantzig-type estimator
$\min_{\rho\in\mathrm{R}^{p}}\|\rho\|_{1}\text{ subject to }\|\widehat{G}\rho-\widehat{M}\|_{\infty}\leq\lambda_{n},$ (7.29)
where $\lambda_{n}$ is an appropriately chosen sequence.
@@ Appendix A Correction Term for Expected Per-Period Utility
Suppose utilities take a linear index form
$u(x,1)=D_{1}(x)^{\prime}\theta_{11},\quad u(x,0)=D_{0}(x)^{\prime}\theta_{10},$ (A.3)
where $\theta_{0}=(\theta_{11},\theta_{10})$. We assume that $\theta_{0}$ is identified via a moment equation
@@ Appendix B Proofs for Results referenced in Main Text
$=\dfrac{\mathrm{E}\bigg{[}1\{K=1\}V_{0}(X)\dfrac{\pi^{0}(S)}{\pi^{1}(S)}\bigg{]}}{{\mathrm{P}}(K=1)}=^{(ii)}\mathrm{E}[w_{0}(X)V_{0}(X)]$
@@
for any $t$ and (iii) from the definition of dynamic Riesz representer $\alpha_{0}(X)$ is in (5.1). Rewriting (5.1) and separating $w_{0}(X)$ from the rest of the terms gives
@@ Appendix C Asymptotic Theory
Appendix D First-Stage Rates for Estimators
@@ D.1. General Case Extremum Estimators
Appendix E First-Stage Rates for Neural Network Estimators of Value Function and Dynamic Riesz Representer
In this Section, we derive mean square convergence rates for estimators of the value function and dynamic Riesz representer based on Theorem D.1. Let $\ell_{V}(z,V,\omega)$ denote an $M$-estimator loss function as in (7.8).
@@ Appendix E First-Stage Rates for Neural Network Estimators of Value Function and Dynamic Riesz Representer
Assumption E.1 is a basic consistency condition for the plug-in estimators of per-period utility $\zeta$. For the dynamic binary choice in Section A, the per-period utility $\zeta(x)$ is a known smooth transformation of the conditional choice probability and a structural parameter given in (2.4)–(A.3).
@@ Appendix F Extension to Nonlinear Models
Let $\alpha_{0}$ be the dynamic Riesz representer corresponding to $w_{D0}$ in (F.4). The following extremum characterization applies for the $M$-estimator loss function
@@ Appendix G Dynamic binary choice revisited
for linear index utilities as in (A.3). Introducing expectation functions for the choices $J=1$ and $J=0$
@@
is a special case of a linear functional (F.8) in Remark F.1 with $\rho(Z)=\alpha_{2}(X,J)-\alpha_{1}(X,J)$. Thus, the true value of $\alpha$ is the dynamic Riesz representer based on weighting function
@@
1:  Partition the set of data indices $\{1,2,\dots,n\}$ into $L$ disjoint subsets of about equal size with where $L$ is an odd number $L\geq 3$. 2:  Let $I^{c}_{\ell}=(Z_{i})_{i\notin I_{\ell}}$ denote the set observations not in $I_{\ell}$. Partition $I^{c}_{\ell}=I^{1}_{\ell}\sqcup I^{2}_{\ell}$ into two halves. For each nuisance parameter $\omega$, let $\widehat{\omega}^{1}_{\ell},\omega^{2}_{\ell}$ denote the estimator computed on $I^{1}_{\ell}$ and $I^{2}_{\ell}$, respectively. Estimate a The choice probability function $p_{0}(x)={\mathrm{P}}(J=1\mid X=x)$ b The expectation functions $V\rightarrow\gamma^{1}(x,V)$ and $V\rightarrow\gamma^{2}(x,V)$ as in (G.8) and (G.7). c The forward operator $\mathrm{A}$ in (7.1) and backward operator $\mathrm{A}^{*}$ in (7.15) 3:  Estimate “slope” and “intercept” value functions by minimizing sample cross-fit objective function $\widehat{V}^{S}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ $\widehat{V}^{I}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ where $\mathcal{V}_{n}$ is some set of functions and $\ell^{S}$ and $\ell^{I}$ are special cases of $M$-estimator losses in (7.8) based on $\zeta^{S}(x)$ and $\zeta^{I}(x)$, respectively. 4:  Estimate dynamic Riesz representer by minimizing sample cross-fit objective function $\widehat{\alpha}_{\ell}=\arg\min_{\alpha\in\mathcal{A}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{2*}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{1*}_{\ell})\bigg{]},$ where $\mathcal{A}_{n}$ is some set of functions and $\ell^{D}$ is a special case of $M$-estimator losses in (F.5) based on $D(Z,V)$. 5:  Estimate the nuisance parameters $(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell})$ using all observations in $I^{c}_{\ell}$. Define a collection of estimated functions as $\widehat{h}_{\ell}=(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{(}V^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{V}^{I}_{\ell},\widehat{(}\zeta^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{\zeta}^{I}_{\ell})$ where $\widehat{\theta}_{\ell}$ is a preliminary estimator of $\theta$ based on a non-orthogonal moment condition. 6:  Estimate the structural parameter using debiased GMM estimator $\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi(Z_{i},\widehat{\theta},\widehat{h}_{\ell})=0$

### v5

@@ 5.1. Dynamic Dual Representation
Remark 5.1 (Implications for Rust (1987) and Hotz and Miller (1993)).
Consider a dynamic binary choice model where one of the actions has a terminal property similar to Rust (1987). Furthermore, suppose the deterministic utilities take a linear index form
$u(x,1)=D_{1}(x)^{\prime}\theta_{11},\quad u(x,0)=D_{0}(x)^{\prime}\theta_{10},$ (5.8)
where $\theta_{0}=(\theta_{11},\theta_{10})$. This parameter can be identified via a semiparametric moment condition whose only nuisance parameter is conditional choice probability Hotz and Miller (1993); Chernozhukov et al. (2022). Stacking this moment condition with (5.5) in Example 2.1 gives a semiparametric moment condition for $(\theta_{0},\delta_{0})$ where neither value function nor any other fixed point of Bellman equation needs to be estimated.
Remark 5.2 (Implications for dynamic discrete choice models as in Aguirregabiria and Mira (2002)).
In a broad class of dynamic discrete choice models, including the one in Duflo et al. (2012), neither action has a terminal choice property. In that case, the structural parameter can be identified by a PMLE moment condition described in Aguirregabiria and Mira (2002) whose debiased analog is proposed in Adusumilli and Eckardt (2019). When the state space is continuously supported, the nuisance components includes choice-specific value functions that are fixed points of a nonparametric IV problem. Appendix G describes a related yet different moment condition that we utilize in Section 4.
@@ 6.1. Average Welfare and Related Averages
Algorithm 1 Step 1 partitions the data into $L$ folds to enable sample splitting. Steps 2 and 3 estimate the value function and the dynamic dual representation on each fold by minimizing cross-fit loss functions. Step 4 estimates the per-period utility $\zeta(\cdot)$ and the bias-correction terms using observations not in fold $\ell$. Step 5 computes the welfare estimate $\widehat{\delta}$ as a bias-corrected plug-in estimator. Step 6 computes the standard error of $\widehat{\delta}$ using the empirical variance of the influence function $\psi$, defined in equation (6.6).
@@ 7.4. Discussion and Related Results
In this Section, we discuss related results. Remark 7.3 discusses cross-fitting. Remark 7.4 introduces a neural network estimator of the value function. Remark 7.5 introduces a neural network estimator of the dynamic dual representation. Remark 7.6 describes the automatic property of the dynamic dual criterion function. Remark 7.7 verifies rate conditions for asymptotic theory.
@@ Appendix A Correction Term for Expected Per-Period Utility
We assume that $\theta_{0}$ is identified via a moment equation
@@ Appendix B Proofs for Results referenced in Main Text
$=\dfrac{\mathrm{E}\bigg{[}1\{K=1\}V_{0}(X)\dfrac{\pi^{0}(S)}{\pi^{1}(S)}\bigg{]}}{{\mathrm{P}}(K=1)}$
$=^{(ii)}\dfrac{\mathrm{E}\bigg{[}1\{K=1\}V_{0}(X)\dfrac{{\mathrm{P}}(K=0\mid S)}{{\mathrm{P}}(K=1\mid S)}\bigg{]}}{{\mathrm{P}}(K=0)}=\mathrm{E}[w_{0}(X)V_{0}(X)]$
@@
for any $t$ and (iii) from the definition of dynamic dual representation $\alpha_{0}(X)$ is in (5.1). Rewriting (5.1) and separating $w_{0}(X)$ from the rest of the terms gives
@@ Appendix C Asymptotic Theory
Appendix D First-Stage Rates for Extremum Estimators
@@ D.1. General Case Extremum Estimators
Appendix E First-Stage Rates for Neural Network Estimators of Value Function and Dynamic Dual Representation
In this Section, we derive mean square convergence rates for estimators of the value function and dynamic dual representation based on Theorem D.1. Let $\ell_{V}(z,V,\omega)$ denote an $M$-estimator loss function as in (7.8).
@@ Appendix E First-Stage Rates for Neural Network Estimators of Value Function and Dynamic Dual Representation
Assumption E.1 is a basic consistency condition for the plug-in estimators of per-period utility $\zeta$. For the dynamic binary choice in Section A, the per-period utility $\zeta(x)$ is a known smooth transformation of the conditional choice probability and a structural parameter given in (2.4)–(5.8).
@@ Appendix F Extension to Nonlinear Models
Let $\alpha_{0}$ be the dynamic dual representation corresponding to $w_{D0}$ in (F.4). The following extremum characterization applies for the $M$-estimator loss function
@@ Appendix G Dynamic binary choice revisited
for linear index utilities as in (5.8). Introducing expectation functions for the choices $J=1$ and $J=0$
@@
is a special case of a linear functional (F.8) in Remark F.1 with $\rho(Z)=\alpha_{2}(X,J)-\alpha_{1}(X,J)$. Thus, the true value of $\alpha$ is the dynamic dual representation based on weighting function
@@
1:  Partition the set of data indices $\{1,2,\dots,n\}$ into $L$ disjoint subsets of about equal size with where $L$ is an odd number $L\geq 3$. 2:  Let $I^{c}_{\ell}=(Z_{i})_{i\notin I_{\ell}}$ denote the set observations not in $I_{\ell}$. Partition $I^{c}_{\ell}=I^{1}_{\ell}\sqcup I^{2}_{\ell}$ into two halves. For each nuisance parameter $\omega$, let $\widehat{\omega}^{1}_{\ell},\omega^{2}_{\ell}$ denote the estimator computed on $I^{1}_{\ell}$ and $I^{2}_{\ell}$, respectively. Estimate a The choice probability function $p_{0}(x)={\mathrm{P}}(J=1\mid X=x)$ b The expectation functions $V\rightarrow\gamma^{1}(x,V)$ and $V\rightarrow\gamma^{2}(x,V)$ as in (G.8) and (G.7). c The forward operator $\mathrm{A}$ in (7.1) and backward operator $\mathrm{A}^{*}$ in (7.15) 3:  Estimate “slope” and “intercept” value functions by minimizing sample cross-fit objective function $\widehat{V}^{S}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{S}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ $\widehat{V}^{I}_{\ell}=\arg\min_{V\in\mathcal{V}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{2}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{I}(Z_{i},\widehat{\omega}^{1}_{\ell})\bigg{]},\qquad\widehat{\omega}=(\widehat{\mathrm{A}},\widehat{p})$ where $\mathcal{V}_{n}$ is some set of functions and $\ell^{S}$ and $\ell^{I}$ are special cases of $M$-estimator losses in (7.8) based on $\zeta^{S}(x)$ and $\zeta^{I}(x)$, respectively. 4:  Estimate dynamic dual representation by minimizing sample cross-fit objective function $\widehat{\alpha}_{\ell}=\arg\min_{\alpha\in\mathcal{A}_{n}}\bigg{[}\sum_{i\in I^{c1}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{2*}_{\ell})+\sum_{i\in I^{c2}_{\ell}}\ell^{D}(Z_{i},\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{\mathrm{A}}^{1*}_{\ell})\bigg{]},$ where $\mathcal{A}_{n}$ is some set of functions and $\ell^{D}$ is a special case of $M$-estimator losses in (F.5) based on $D(Z,V)$. 5:  Estimate the nuisance parameters $(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell})$ using all observations in $I^{c}_{\ell}$. Define a collection of estimated functions as $\widehat{h}_{\ell}=(\widehat{p}_{\ell},\widehat{\gamma}^{2}_{\ell},\widehat{\gamma}^{1}_{\ell},\widehat{(}V^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{V}^{I}_{\ell},\widehat{(}\zeta^{S}_{\ell})^{\prime}\widehat{\theta}_{\ell}+\widehat{\zeta}^{I}_{\ell})$ where $\widehat{\theta}_{\ell}$ is a preliminary estimator of $\theta$ based on a non-orthogonal moment condition. 6:  Estimate the structural parameter using debiased GMM estimator $\frac{1}{n}\sum_{\ell=1}^{L}\sum_{i\in I_{\ell}}\psi(Z_{i},\widehat{\theta},\widehat{h}_{\ell})=0$